OpenStudy (anonymous):

mrs. torres is choosing glazes for her pottery. Her choices are shown in the table: Color Texture Polish Blue smooth shiny Red Rough Dull Green What is the probability Mrs. Torres will chose a red, smooth, and shiny glaze? a. 1/8 b. 1/3 c. 3/7 d. 1/12 Please help me with how to begin, someone! Please.

OpenStudy (kropot72):

What is the probability of choosing red out of the 3 colors?

OpenStudy (anonymous):

Probability of choosing red = 1/3 Probability of choosing smooth = 1/2 probability of choosing shiny = 1/2 \[\frac 1 3 \times \frac 12 \times \frac 12 =\frac 1{12} \]

OpenStudy (anonymous):

P(red, smooth, and shiny glaze) = P(red) x P(smooth) x P(shiny) = ....

OpenStudy (anonymous):

ok...so green is not included at all? Thanks for your help!!

OpenStudy (anonymous):

You need a red, not a green :)

OpenStudy (anonymous):

Thank you!!!

OpenStudy (anonymous):

welcome!